# Directly Denoising Diffusion Model

### v1

@@ Abstract
In this paper, we present the Directly Denoising Diffusion Model (DDDM): a simple and generic approach for generating realistic images with few-step sampling, while multistep sampling is still preserved for better performance. DDDMs require no delicately designed samplers nor distillation on pre-trained distillation models. DDDMs train the diffusion model conditioned on an estimated target that was generated from previous training iterations of its own. To generate images, samples generated from the previous time step are also taken into consideration, guiding the generation process iteratively. We further propose Pseudo-LPIPS, a novel metric loss that is more robust to various values of hyperparameter. Despite its simplicity, the proposed approach can achieve strong performance in benchmark datasets. Our model achieves FID scores of 2.57 and 2.33 on CIFAR-10 in one-step and two-step sampling respectively, surpassing those obtained from GANs and distillation-based models. By extending the sampling to 1000 steps, we further reduce FID score to 1.79, aligning with state-of-the-art methods in the literature. For ImageNet 64x64, our approach stands as a competitive contender against leading models.
@@ 2.2 Stochastic Differential Equation Formulation
By replacing $\nabla_{\bm{x}}\log q_{t}\left(\bm{x}_{t}\right)$ with a neural network based estimation $\mathbf{s}_{\bm{\theta}}\left(\mathbf{x}_{t},t\right)$, Song et al. (2020b) obtained the neural ODE:
$\mathrm{d}\mathbf{x}_{t}=-\frac{1}{2}\beta(t)\left[\mathbf{x}_{t}+\mathbf{s}_{\bm{\theta}}\left(\mathbf{x}_{t},t\right)\right]\mathrm{d}t$
@@ 3 Directly Denosing Diffusion Models
By approximate $\mathbf{f}$, we can then recover the original image $\mathbf{x}_{0}$. We define a neural network-parameterized function $\mathbf{f}_{\bm{\theta}}$, which is employed to estimate the solution of the PF ODE and thereby recover the original image state at time 0. The predictive model is represented as:
$\mathbf{f}_{\bm{\theta}}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)=\mathbf{x}_{t}-\mathbf{F}_{\bm{\theta}}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)$ (3)
where $\mathbf{F}_{\bm{\theta}}$ is the neural network function parameterized by the weights $\bm{\theta}$. To achieve a good recovery of the initial state $\mathbf{x}_{0}$, $\mathbf{f}_{\bm{\theta}}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)\approx\mathbf{f}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)$ need to be ensured.
@@ 3.1 Iterative solution
Eq. (3) shows that our neural network $\mathbf{F}_{\mathbf{\theta}}$ requires $\mathbf{x}_{0}$ as input, which is not applicable during sample generation. To unify the training and inference within the same framework, we propose an iterative update rule to estimate the initial state $\mathbf{x}_{0}$ in a dynamic system. This iterative process is formally defined by the following update equation:
$\mathbf{x}_{0}^{(n+1)}=\mathbf{x}_{t}-\mathbf{F}_{\bm{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right)$ (4)
@@
$\mathcal{L}_{\text{DDDM}}^{(n)}(\bm{\theta}):=$ $\mathbb{E}_{t\sim\mathcal{U}[1,T]}[\mathbb{E}_{\mathbf{x}_{0}\sim p_{\text{data }}(\mathbf{x}_{0})}$ (5)
@@
$[d(\mathbf{f}_{\bm{\theta}}(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t),\mathbf{x}_{0})]]]$
@@
Training. Each data sample $\mathbf{x}_{0}$ is chosen randomly from the dataset, following the probability distribution $p_{\text{data }}(\mathbf{x}_{0})$. This initial data point forms the basis for generating a trajectory. Next, we randomly sample a timestep $t\sim\mathcal{U}[1,T]$, and obtain its noisy variant $\mathbf{x}_{t}$ from distribution $\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right)$. we play the reparameterization trick to rewrite $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. For current training epoch $n$, our model takes noisy data $\textbf{x}_{t}$ and timestep $t$, as well as the corresponding estimated target from previous epoch $\textbf{x}_{0}^{(n-1)}$ as inputs, predicts a new approximation $\textbf{x}_{0}^{(n)}$, which will be utilized in the next training epoch for the same target sample. DDDM is trained by minimizing the loss following Eq. 5. The full procedure of training DDDM is summarized in Algorithm 1.
Sampling. The generation of samples is facilitated through the use of a well-trained DDDM, denoted as $\bm{f}_{\bm{\theta}}(\cdot,\cdot)$. The process begins by drawing from the initial Gaussian distribution, where both $\mathbf{x}_{0}^{(0)}$ and $\mathbf{x}_{T}$ are sampled from $\mathcal{N}\left(\mathbf{0},\bm{I}\right)$. Subsequently, these noise vectors and embedding of $T$ are passed through the DDDM model to obtain $\mathbf{x}_{0}^{\text{est}}=\mathbf{f}_{\bm{\theta}}\left(\mathbf{x}_{0}^{(0)},\mathbf{x}_{T},T\right)$. This approach is noteworthy for its efficiency, as it requires only a single forward pass through the model. Our model also supports a multistep sampling procedure for enhanced sample quality. Detail can be found in Algorithm 2.
@@
  Input: image dataset $D,$ $T$, model parameter $\bm{\theta}$ initialize $\mathbf{x}_{0}^{(0)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, epoch $n\leftarrow 0$
@@
     Sample $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
     $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-{\bar{\alpha}}_{t}}\bm{\epsilon}$
     $\mathbf{x}_{0}^{(n+1)}\leftarrow\mathbf{x}_{t}-\mathbf{F}_{\bm{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right)$
     $\mathcal{L}_{\text{DDDM}}^{(n+1)}(\bm{\theta})\leftarrow d\left(\bm{f}_{\bm{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right),\mathbf{x}_{0}\right)$
     $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}\mathcal{L}\left(\bm{\theta}\right)$
@@
  Input: $T$, trained model parameter $\bm{\theta}$, sampling step $s$ $\mathbf{x}_{0}^{(0)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
@@
     $\mathbf{x}_{0}^{(n+1)}\leftarrow\mathbf{x}_{T}-\mathbf{F}_{\bm{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{T},T\right)$
@@
If the loss function $\mathcal{L}_{\text{DDDM}}^{(n)}\left(\bm{\theta}\right)\rightarrow 0$ as $n\rightarrow\infty$, it can be shown that as $n\rightarrow\infty$,
$\sup_{\mathbf{x}_{0}}\left(\mathbf{f}_{\bm{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right)-\mathbf{f}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)\right)\rightarrow 0$ (6)
@@
As $n\rightarrow\infty,\mathcal{L}_{\text{DDDM}}^{(n)}\left(\bm{\theta}\right)\rightarrow 0$, we have
$\mathbb{E}\left[d\left(\mathbf{f}_{\bm{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right),\mathbf{x}_{0}\right)\right]\rightarrow 0$
@@
$d\left(\mathbf{f}_{\bm{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right),\mathbf{x}_{0}\right)\rightarrow 0$
@@
$\mathbf{f}_{\bm{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right)\rightarrow\mathbf{x}_{0}\text{, as }n\rightarrow\infty,$
@@
$\sup_{\mathbf{x}_{0}}d\left(\mathbf{f}_{\bm{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right)-\mathbf{f}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)\right)\rightarrow 0.$
@@
(i) For all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{D}$, time $t\in\mathcal{U}[1,T]$, $\mathbf{x}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right)$, $\mathbf{y}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right)$, the function $\mathbf{f}_{\bm{\theta}}$ satisfies the uniform Lipschitz condition:
$\sup_{\bm{\theta}}\left\|\mathbf{f}_{\bm{\theta}}\left(\mathbf{x},\mathbf{x}_{t},t\right)-\mathbf{f}_{\bm{\theta}}\left(\mathbf{y},\mathbf{y}_{t},t\right)\right\|_{2}\leq L\|\mathbf{x}-\mathbf{y}\|_{2}$
@@
(ii) There exists a non-negative function $L(\mathbf{x})$ such that for all $\mathbf{x}\in\mathbb{R}^{D}$, time $t\in\mathcal{U}[1,T]$ and $\mathbf{x}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right)$, the function $\mathbf{f}_{\bm{\theta}}$ is uniformly bounded in $\bm{\theta}$ :
$\sup_{\bm{\theta}}\left\|\mathbf{f}_{\bm{\theta}}\left(\mathbf{x},\mathbf{x}_{t},t\right)\right\|_{2}\leq L(\mathbf{x})<\infty\text{. }$
If there exists a $\bm{\tilde{{\theta}}}$ such that the loss function $\mathcal{L}_{\text{DDDM}}\left(\bm{\tilde{{\theta}}}\right)\rightarrow 0$ as the iteration number $n$ becomes sufficiently large. Let $p_{\bm{\tilde{{\theta}}}}(\cdot)$ denote the pushforward distribution of $p_{T}$ induced by $\mathbf{f}_{\bm{\tilde{{\theta}}}}(\cdot,\mathbf{x}_{T},T)$.
Then, under these conditions, the discrepancy between pushforward distribution $p_{\bm{\tilde{{\theta}}}}(\cdot)$ and the data distribution $p_{\text{data }}(\cdot)$ measured in the uniform norm, converges to 00 as the number of iterations $n$ approaches infinity: $\left\|p_{\bm{\tilde{{\theta}}}}(\cdot)-p_{\text{data }}(\cdot)\right\|_{\infty}\rightarrow 0$ as iteration $n\rightarrow\infty$.
@@
when $n$ is large enough. Then we conclude that the pushforward distribution of $\mathbf{x}_{T}$, say $p_{\bm{\tilde{\theta}}}(\cdot)$, converges in distribution to the data distribution $p_{\text{data}(\cdot)}$. Since for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{D},t\sim\mathcal{U}[1,T]$, and $\bm{\theta}$,$\left\{\mathbf{f}_{\bm{\theta}}\right\}_{\bm{\theta}}$ is uniform Lipschitz $\left\|\mathbf{f}_{\bm{\theta}}\left(\mathbf{x},\mathbf{x}_{t},t\right)-\mathbf{f}_{\bm{\theta}}\left(\mathbf{y},\mathbf{y}_{t},t\right)\right\|_{2}\leq L\left\|\mathbf{x}-\mathbf{y}\right\|_{2},\quad$$\left\{\mathbf{f}_{\bm{\theta}}\right\}_{\bm{\theta}}$ is asymptotically uniformly equicontinuous (a.u.e.c.). Additionally, $\left\{\mathbf{f}_{\bm{\theta}}\right\}_{\bm{\theta}}$ is uniform bounded in $\bm{\tilde{\theta}}$. Thus, by the converse of Scheffé’s theorem (Boos, 1985; Sweeting, 1986), it suggests $\left\|p_{\bm{\tilde{\theta}}}(\cdot)-p_{\text{data }}(\cdot)\right\|_{\infty}\rightarrow 0$ as $n$ sufficiently large. ∎

### v2

@@ Abstract
In this paper, we present the Directly Denoising Diffusion Model (DDDM): a simple and generic approach for generating realistic images with few-step sampling, while multistep sampling is still preserved for better performance. DDDMs require no delicately designed samplers nor distillation on pre-trained distillation models. DDDMs train the diffusion model conditioned on an estimated target that was generated from previous training iterations of its own. To generate images, samples generated from the previous time step are also taken into consideration, guiding the generation process iteratively. We further propose Pseudo-LPIPS, a novel metric loss that is more robust to various values of hyperparameter. Despite its simplicity, the proposed approach can achieve strong performance in benchmark datasets. Our model achieves FID scores of 2.57 and 2.33 on CIFAR-10 in one-step and two-step sampling respectively, surpassing those obtained from GANs and distillation-based models. By extending the sampling to 1000 steps, we further reduce FID score to 1.79, aligning with state-of-the-art methods in the literature. Our code is available at https://github.com/TheLuoFengLab/DDDM.
@@ 2.2 Stochastic Differential Equation Formulation
By replacing $\nabla_{\boldsymbol{x}}\log q_{t}\left(\boldsymbol{x}_{t}\right)$ with a neural network based estimation $\mathbf{s}_{\boldsymbol{\theta}}\left(\mathbf{x}_{t},t\right)$, Song et al. (2020b) obtained the neural ODE:
$\mathrm{d}\mathbf{x}_{t}=-\frac{1}{2}\beta(t)\left[\mathbf{x}_{t}+\mathbf{s}_{\boldsymbol{\theta}}\left(\mathbf{x}_{t},t\right)\right]\mathrm{d}t$
@@ 3 Directly Denosing Diffusion Models
By approximate $\mathbf{f}$, we can then recover the original image $\mathbf{x}_{0}$. We define a neural network-parameterized function $\mathbf{f}_{\boldsymbol{\theta}}$, which is employed to estimate the solution of the PF ODE and thereby recover the original image state at time 0. The predictive model is represented as:
$\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)=\mathbf{x}_{t}-\mathbf{F}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)$ (3)
where $\mathbf{F}_{\boldsymbol{\theta}}$ is the neural network function parameterized by the weights $\boldsymbol{\theta}$. To achieve a good recovery of the initial state $\mathbf{x}_{0}$, $\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)\approx\mathbf{f}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)$ need to be ensured.
@@ 3.1 Iterative solution
Eq. (3) shows that our neural network $\mathbf{F}_{\boldsymbol{\theta}}$ requires $\mathbf{x}_{0}$ as input, which is not applicable during sample generation. To unify the training and inference within the same framework, we propose an iterative update rule to estimate the initial state $\mathbf{x}_{0}$ in a dynamic system. This iterative process is formally defined by the following update equation:
$\mathbf{x}_{0}^{(n+1)}=\mathbf{x}_{t}-\mathbf{F}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right)$ (4)
@@
$\mathcal{L}_{\text{DDDM}}^{(n)}(\boldsymbol{\theta}):=$ $\mathbb{E}_{t\sim\mathcal{U}[1,T]}[\mathbb{E}_{\mathbf{x}_{0}\sim p_{\text{data }}(\mathbf{x}_{0})}$ (5)
@@
$[d(\mathbf{f}_{\boldsymbol{\theta}}(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t),\mathbf{x}_{0})]]]$
@@
Training. Each data sample $\mathbf{x}_{0}$ is chosen randomly from the dataset, following the probability distribution $p_{\text{data }}(\mathbf{x}_{0})$. This initial data point forms the basis for generating a trajectory. Next, we randomly sample a timestep $t\sim\mathcal{U}[1,T]$, and obtain its noisy variant $\mathbf{x}_{t}$ from distribution $\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right)$. we play the reparameterization trick to rewrite $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. For current training epoch $n$, our model takes noisy data $\textbf{x}_{t}$ and timestep $t$, as well as the corresponding estimated target from previous epoch $\textbf{x}_{0}^{(n-1)}$ as inputs, predicts a new approximation $\textbf{x}_{0}^{(n)}$, which will be utilized in the next training epoch for the same target sample. DDDM is trained by minimizing the loss following Eq. 5. The full procedure of training DDDM is summarized in Algorithm 1.
Sampling. The generation of samples is facilitated through the use of a well-trained DDDM, denoted as $\boldsymbol{f}_{\boldsymbol{\theta}}(\cdot,\cdot)$. The process begins by drawing from the initial Gaussian distribution, where both $\mathbf{x}_{0}^{(0)}$ and $\mathbf{x}_{T}$ are sampled from $\mathcal{N}\left(\mathbf{0},\boldsymbol{I}\right)$. Subsequently, these noise vectors and embedding of $T$ are passed through the DDDM model to obtain $\mathbf{x}_{0}^{\text{est}}=\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0}^{(0)},\mathbf{x}_{T},T\right)$. This approach is noteworthy for its efficiency, as it requires only a single forward pass through the model. Our model also supports a multistep sampling procedure for enhanced sample quality. Detail can be found in Algorithm 2.
@@
  Input: image dataset $D,$ $T$, model parameter $\boldsymbol{\theta}$ initialize $\mathbf{x}_{0}^{(0)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, epoch $n\leftarrow 0$
@@
     Sample $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
     $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-{\bar{\alpha}}_{t}}\boldsymbol{\epsilon}$
     $\mathbf{x}_{0}^{(n+1)}\leftarrow\mathbf{x}_{t}-\mathbf{F}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right)$
     $\mathcal{L}_{\text{DDDM}}^{(n+1)}(\boldsymbol{\theta})\leftarrow d\left(\boldsymbol{f}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right),\mathbf{x}_{0}\right)$
     $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}\left(\boldsymbol{\theta}\right)$
@@
  Input: $T$, trained model parameter $\boldsymbol{\theta}$, sampling step $s$ $\mathbf{x}_{0}^{(0)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
@@
     $\mathbf{x}_{0}^{(n+1)}\leftarrow\mathbf{x}_{T}-\mathbf{F}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{T},T\right)$
@@
If the loss function $\mathcal{L}_{\text{DDDM}}^{(n)}\left(\boldsymbol{\theta}\right)\rightarrow 0$ as $n\rightarrow\infty$, it can be shown that as $n\rightarrow\infty$,
$\sup_{\mathbf{x}_{0}}\left(\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right)-\mathbf{f}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)\right)\rightarrow 0$ (6)
@@
As $n\rightarrow\infty,\mathcal{L}_{\text{DDDM}}^{(n)}\left(\boldsymbol{\theta}\right)\rightarrow 0$, we have
$\mathbb{E}\left[d\left(\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right),\mathbf{x}_{0}\right)\right]\rightarrow 0$
@@
$d\left(\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right),\mathbf{x}_{0}\right)\rightarrow 0$
@@
$\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right)\rightarrow\mathbf{x}_{0}\text{, as }n\rightarrow\infty,$
@@
$\sup_{\mathbf{x}_{0}}d\left(\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x}_{0}^{(n)},\mathbf{x}_{t},t\right)-\mathbf{f}\left(\mathbf{x}_{0},\mathbf{x}_{t},t\right)\right)\rightarrow 0.$
@@
(i) For all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{D}$, time $t\in\mathcal{U}[1,T]$, $\mathbf{x}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right)$, $\mathbf{y}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right)$, the function $\mathbf{f}_{\boldsymbol{\theta}}$ satisfies the uniform Lipschitz condition:
$\sup_{\boldsymbol{\theta}}\left\|\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x},\mathbf{x}_{t},t\right)-\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{y},\mathbf{y}_{t},t\right)\right\|_{2}\leq L\|\mathbf{x}-\mathbf{y}\|_{2}$
@@
(ii) There exists a non-negative function $L(\mathbf{x})$ such that for all $\mathbf{x}\in\mathbb{R}^{D}$, time $t\in\mathcal{U}[1,T]$ and $\mathbf{x}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{x},\left(1-\bar{\alpha}_{t}\right)\mathbf{I}\right)$, the function $\mathbf{f}_{\boldsymbol{\theta}}$ is uniformly bounded in $\boldsymbol{\theta}$ :
$\sup_{\boldsymbol{\theta}}\left\|\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x},\mathbf{x}_{t},t\right)\right\|_{2}\leq L(\mathbf{x})<\infty\text{. }$
If there exists a $\boldsymbol{\tilde{{\theta}}}$ such that the loss function $\mathcal{L}_{\text{DDDM}}\left(\boldsymbol{\tilde{{\theta}}}\right)\rightarrow 0$ as the iteration number $n$ becomes sufficiently large. Let $p_{\boldsymbol{\tilde{{\theta}}}}(\cdot)$ denote the pushforward distribution of $p_{T}$ induced by $\mathbf{f}_{\boldsymbol{\tilde{{\theta}}}}(\cdot,\mathbf{x}_{T},T)$.
Then, under these conditions, the discrepancy between pushforward distribution $p_{\boldsymbol{\tilde{{\theta}}}}(\cdot)$ and the data distribution $p_{\text{data }}(\cdot)$ measured in the uniform norm, converges to 00 as the number of iterations $n$ approaches infinity: $\left\|p_{\boldsymbol{\tilde{{\theta}}}}(\cdot)-p_{\text{data }}(\cdot)\right\|_{\infty}\rightarrow 0$ as iteration $n\rightarrow\infty$.
@@
when $n$ is large enough. Then we conclude that the pushforward distribution of $\mathbf{x}_{T}$, say $p_{\boldsymbol{\tilde{\theta}}}(\cdot)$, converges in distribution to the data distribution $p_{\text{data}(\cdot)}$. Since for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{D},t\sim\mathcal{U}[1,T]$, and $\boldsymbol{\theta}$,$\left\{\mathbf{f}_{\boldsymbol{\theta}}\right\}_{\boldsymbol{\theta}}$ is uniform Lipschitz $\left\|\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{x},\mathbf{x}_{t},t\right)-\mathbf{f}_{\boldsymbol{\theta}}\left(\mathbf{y},\mathbf{y}_{t},t\right)\right\|_{2}\leq L\left\|\mathbf{x}-\mathbf{y}\right\|_{2},\quad$$\left\{\mathbf{f}_{\boldsymbol{\theta}}\right\}_{\boldsymbol{\theta}}$ is asymptotically uniformly equicontinuous (a.u.e.c.). Additionally, $\left\{\mathbf{f}_{\boldsymbol{\theta}}\right\}_{\boldsymbol{\theta}}$ is uniform bounded in $\boldsymbol{\tilde{\theta}}$. Thus, by the converse of Scheffé’s theorem (Boos, 1985; Sweeting, 1986), it suggests $\left\|p_{\boldsymbol{\tilde{\theta}}}(\cdot)-p_{\text{data }}(\cdot)\right\|_{\infty}\rightarrow 0$ as $n$ sufficiently large. ∎